OpenStudy (anonymous):

Write an equation of the line that passes through the given point and is perpendicular to the given equation : (-3,-2); 2x+4y=7 and (4,-3); y= 3/7x-5

OpenStudy (anonymous):

Please help!

OpenStudy (zehanz):

Normal vector of 2x+4y=7 is \[\left(\begin{matrix}2 \\ 4\end{matrix}\right)\]Then a normal vector of lines perpendicular to it is:\[\left(\begin{matrix}-4 \\ 2\end{matrix}\right)\]leading to the equation -4x+2y=c. If you substitute x=-3 and y=-2 in this equation, you get c.

OpenStudy (anonymous):

1. The slope for the new equation would be 2x becaus eof the recipical Y-Y1= m(X-X1) Y-(-2)=2(X+3) y+2=2x+6 y=2x+4 2. 7/3x is the new slope Y-Y1= m(x-x1) Y-(-3)=-7/3(x-4) Y+3=-7/3x+9 1/3 y=-7/3x+6 1/3

OpenStudy (anonymous):

^ its wrong

OpenStudy (zehanz):

x=-3, y=-2: -4*-3+2*-2=12-4=8, so eq is -4x+2y=8, or (simpler) -2x+y=4